OpenStudy (clara1223):

Calculate the derivative of the given function f(x)=8sin^3(√x) a) f′(x)=24sin^2(√x)cos(√x) b) f′(x)=48sin^2(√x)cos(√x)/x√x c) f′(x)=24cos(√x) d) f′(x)=12sin^2(√x)cos(√x)/√x e) f′(x)=12cos(√x)/√x

OpenStudy (clara1223):

yes, one second

OpenStudy (jhannybean):

Sorry, I made a typo. \[f(x) = 8\sin^3(x^{1/2})\]

OpenStudy (clara1223):

actually could you help me with this really quickly? its more pressing: Evaluate (g∘f)′(6), given that f(4)=4, f(5)=6, f(6)=6, g(4)=5, g(5)=5, g(6)=4, f′(4)=6, f′(5)=5, f′(6)=4, g′(4)=5, g′(5)=5, g′(6)=4 a) 16 b) 19 c) 18 d) 17 e) 15

OpenStudy (clara1223):

sorry, thanks for all your help, im just on a time crunch.

OpenStudy (jhannybean):

You're not taking a test are you?

OpenStudy (clara1223):

no, I just have to be somewhere soon

OpenStudy (jhannybean):

Oh, alright.

OpenStudy (clara1223):

And this homework is due tomorrow

OpenStudy (jhannybean):

So i assume : \[(g~\circ ~ f)'(x) \qquad \implies \qquad \frac{d}{dx}(g(f(x)) = g'(f(x)) \cdot f'(x)\]

OpenStudy (jhannybean):

where \(x=6\)

OpenStudy (jhannybean):

Ack, I've got to head off for a little while, I will be back very soon.

OpenStudy (jhannybean):

@DanJS @jim_thompson5910 mind taking over? Thanks!

OpenStudy (danjs):

i am late, are you still working on these things?

OpenStudy (danjs):

yes, that one would be just the chain rule, and they list the values you may need for the functions and their derivatives \[\qquad \frac{d}{dx}(g(f(6)) = g'(f(6)) \cdot f'(6) \]

OpenStudy (danjs):

f(6) = 6 f '(6) = 4 g'(6)=4 so \[\qquad \frac{d}{dx}(g(f(x)) = g'(f(x)) \cdot f'(x) = g'(6)*f'(6) = 4*4\]

OpenStudy (danjs):

To find f'(x) you use the chain rule... \[\frac{ d }{ dx }f(x) = \frac{ d }{ du }*\frac{ du }{ dx } f(x)\]

OpenStudy (danjs):

If you have a function inside a function , like f(g(x)) f(x) = sin(x^3) I usually think something like; " f ' (x) is equal to derivative of sin(inside function left alone) times the derivative of the inside function" f ' (x) = cos(x^3) * 2x^2 thats it, goodluck

OpenStudy (danjs):

similar with the given function f(x) = 8*sin(x^(1/2))^3 start with outside and leave inside function same, except here you have to use the rule two times, u^3, sin(u) and x^(1/2) start from outside power function \[f ' (x) = 3*8*[\sin(x^{1/2})]^2~*~\cos(x^{1/2})*\frac{ 1 }{ 2 }*x^{-1/2}\]

OpenStudy (danjs):

simplifies to D i think ... check it out,

OpenStudy (danjs):

keep leaving the inner functions alone, and work outside to inside, power rule - derivative of sin() - derivative of root x

OpenStudy (danjs):

\[[f(g(h(x)))] ' = f'(g(h(x))) *g'(h(x)) * h'(x)\]

OpenStudy (danjs):

you can have as many nested functions as you want, same idea